OpenStudy (anonymous):

What is the simplified form of the expression? (m^3/4c^5)^4

OpenStudy (anonymous):

m^12/256c^20

OpenStudy (anonymous):

sai that wasn't one of my answers ;/

OpenStudy (anonymous):

THis involves using two the law of exponents. \[(a^{b})^{c} = a^{bc}\]\[(ab)^{c} = a^{c}b^c\] \[(\frac{m^{3}}{4c^{5}})^{4} = \frac{m^{3 \times 4}}{4^{4}{c^{5 \times 4}}} = \frac{m^{12}}{256c^{20}}\]I got the same thing.

OpenStudy (anonymous):

would it still be right if it was written backwards? the closest i got was 256c^20/m^12

OpenStudy (anonymous):

no that is wrong, it means you've reciprocated the answer which you shouldn't have: this is the solution: (m^3/4c^5)^4 raise each element to the power 4 hence: m^3*4 ; 4^4 ; and C ^5*4 m^12 ;256 and c^20 the answer will then be: m^12/ 256 c^20

OpenStudy (anonymous):

no i ended up getting it right